### v1

@@ 5.2.2 Spatial + Temporal
Many models use a two-step approach, where the spatial domain is first processed by a spatial perception module to extract spatial features. These features are then passed into a sequential model for temporal aggregation, resulting in a combined spatial-temporal representation of the input sequence. For instance, GCN-BERT [176] utilizes a two-layer GCN for spatial modeling, followed by BERT for temporal modeling.
Other general-purpose methods adopt a multi-stream, two-stage pose-based approach. In [129], three streams, each containing three 1D convolutional layers, are combined with a parallel module of average pooling layers. The extracted features are then concatenated and input into an LSTM, followed by a fully connected layer and softmax for classification. Similarly, STr-GCN [165] employs a two-stage model where a GCN spatial feature extractor is first applied to obtain intra-frame features and then passes the features through a transformer encoder for modeling inter-frame temporal relationships. The final classification scores are generated by an MLP.
@@ 5.2.3 Short-Term Spatiotemporal + Long-Term Temporal
For example, gesture recognition in autonomous driving has become an important research area. In [59], the authors propose a pose-based approach called MG-GCT for recognizing police traffic gestures. The model uses two graph convolution streams: one to process skeleton joints and another to analyze joint motions. The features extracted from both streams are fused and passed through a transformer encoder for temporal modeling, with the final classification performed by an MLP. In another approach, TD-GCN [107] introduces a pose-based method for IDHGR. Unlike traditional methods that use fixed skeleton topologies, TD-GCN employs temporal-sensitive topology learning to enhance feature quality. First, graph convolution is used to extract high-level spatiotemporal representations, which are then used to calculate temporal and channel-independent adjacency matrices. These matrices help model the skeleton, and the features are fused using the proposed TCF module.
@@ 5.2.4 Temporal
In [139], the authors proposed a pose-based method for surgical robot teleoperation. Preprocessed skeleton joints are fed into an LSTM, and the resulting features are passed to an MLP for final classification. The model was evaluated on a custom dataset. Similarly, in [151], a cross-domain gesture recognition approach was developed using geometric features (e.g., bone angles), which are processed by a TCN. The resulting features are summarized into a sequence descriptor, and a linear classifier predicts the final class.
To address data scarcity, the authors in [128] introduced AVSN, an adversarial learning-based data augmentation framework. The method jointly optimizes augmentation parameters and the HGR classifier. A BiLSTM serves as the discriminator, while an MLP acts as the generator. The generator produces challenging augmented data that exploits weaknesses in the discriminator, which in turn learns to correctly classify gestures.
@@ 5.2.5 Other
Some architectures adopt more complex designs or alternative skeleton representations, placing them outside the previously defined categories.
In [1], the authors combined a Fourier CNN with a BiLSTM to extract spatiotemporal features from 3D skeleton joints and perform ISLR. A trainable binary mask was also used for feature selection to eliminate redundancy. Similarly, authors in [10] performed ISLR using a feature thresholding technique that transforms skeleton features based on their variance and importance. These transformed features undergo exhaustive search for selection and are then processed by a deep CNN for classification.
To address limitations of color-coded feature maps, the authors in [89] proposed the Joint Motion Affinity Map (JMAM) as a new skeleton representation for ISLR. JMAM is input to a custom multi-stream attention-based 2DCNN, which extracts features and performs classification. Key frame selection was explored in [46] through the Frame Selection Network (FS-Net), trained using reinforcement learning. The agent selects frames for classification by DD-Net and receives a reward of +1 or -1 based on the correctness of the classification outcome. Finally, [101] proposed a lightweight dual-stream method to fully capture spatiotemporal features. One stream employs stacked attention-based modules, while the other uses an autoencoder. The outputs of both are classified independently, and their predictions are averaged to yield the final decision.
@@ 5.3.1 General Purpose
In [100], a dual-stream 3DCNN followed by average fusion was proposed for RGB-D gesture recognition. To filter out redundant visual features ($\upsilon$) and retain only the most discriminative components ($z$), a memory bank mechanism was introduced. The training objective maximizes mutual information between $z$ and the gesture class $y$, while minimizing that between $z$ and $\upsilon$. The memory bank is updated via read/write operations during training. Similarly, addressing redundancy, authors in [206] applied Neural Architecture Search (NAS) and attention mechanisms. A modified I3D backbone replaces some inception modules with learnable NAS cells. These progressively transition from small-kernel to dilated convolutions. Dynamic Static Attention (DSA) modules focus on motion and hand/arm regions, aided by a parallel HeatmapNet trained jointly through an auxiliary loss. For RGB-D input, late fusion is used.
To move beyond the limitations of late fusion and exploit modality interactions more effectively, [207] proposed a decoupling-recoupling framework. Independent spatial (with a Focused Region Proposal (FRP) module) and temporal modeling (via attention) are first performed to extract modality-specific features. These are then fused mid-way using a shared MLP-AE module. Multi-level fusion was further studied in [60], where 3DCNN backbones are combined with a Convolutional Transformer Fusion Block (CTFB). Injected at multiple layers, CTFB fuses intermediate features across streams for better spatial-temporal integration.
To address data scarcity, several works explored zero-shot learning (ZSL) by incorporating textual gesture descriptions encoded by pretrained language models (mainly BERT). In [146], a dual-stream model combined a ViT for spatial modeling and an LSTM-MLP for temporal modeling, learning visual features aligned with textual embeddings. Similarly, [147] fused skeleton-derived geometric features with RGB-C3D-LSTM-AE representations into a shared semantic space for classification. To reduce computational cost, [53] replaced the transformer’s self-attention mechanism with convolutional layers and a gated feedforward network, resulting in the lightweight ConvMixFormer architecture.
In [180], knowledge distillation and contrastive learning were employed to enhance multimodal performance. Three 3DCNNs (for RGB, Depth, and Optical Flow) extract features fused via Coordinate Attention (CA). I3D serves as the teacher, and ResNet as the student. CLIP is also integrated for zero-shot gesture representation. The limitations of standard 3DCNNs were explored in [114], where 3D Central Difference Convolution (3D-CDC) replaces standard 3D convolutions. Residual hybrid blocks and a transformer module handle temporal modeling. [195] also adopted 3D-CDC in a multi-stream, multi-rate setup with multi-level fusion. NAS optimizes the configuration of CDC types, inter-stream connections, and frame rates.
@@ 5.3.2 Sign Language Understanding
Several recent works have proposed multimodal and pose-centric architectures for Isolated Sign Language Recognition (ISLR), focusing on combining RGB and skeleton data.
@@
A prominent line of research is the SignBERT framework [68], which uses a spatial GCN with temporal and hand chirality encoding to extract gesture features. These are fed into a transformer encoder, while a hand-model-aware decoder reconstructs masked inputs during self-supervised pretraining. Predictions from a parallel RGB network are fused at inference time. An enhanced version, SignBERT+ [71], introduces spatial-temporal position encoding and an improved fusion module for RGB and pose data. The architecture is illustrated in Fig. 10. Similarly, MASA [203] adopts the same GCN+Transformer backbone, trained with contrastive learning on pose reconstruction. BEST [202] extends this by embedding pose triplets (left hand, right hand, body) with BERT and applying late fusion with RGB data.
@@
Figure 10: The pose-based stream of the SignBERT+ [71], an extension of the SignBERT [68], during SSL pretraining. A GCN is first utilized to extract the gesture features, which are after combined with spatial-temporal positing encoding and subsequently pass through a transformer encoder. During pretraining, a MANO-based decoder is placed after the transformer to perform reconstruction of the masked input. The decoder is replaced with a prediction head during downstream fine-tuning.
@@
To improve feature invariance, a contrastive approach was proposed in [204], enforcing consistency among hand, trunk, and motion features. The same GCN+Transformer architecture is used, with late fusion of pose and RGB predictions. The SAM-SLR framework [82] combines six streams (RGB, depth, skeleton) with late fusion of predictions. SAM-SLR-v2 [81] adds a 3D hand graph stream and learnable fusion weights (Fig. 11).
Figure 11: The architecture of SAM-SLR-v2 [81], which is based on the SAM-SLR [82]. The architecture consists of seven streams, regarding various modalities, including RGB frames, optical flow, pose heatmaps and skeleton joints. The predictions of each branch are fused in a learnable late manner.
In VTN-HC/PF [36], hand-cropped images are processed via a 2D-CNN for spatial modeling, followed by a transformer and MLP for temporal and final predictions. The PF variant incorporates geometrical features from pose data. Motion History Images (MHIs) were employed in [119], where a 3DCNN with attention and residual connections analyzes MHIs, while a dual-stream version adds a 2DCNN for RGB data, with predictions fused via weighted summation.
To address data scarcity, ZSSLR [216] adopts a CLIP-based zero-shot framework using RGB and pose data. In [24], a few-shot cross-lingual method is proposed, combining two 3DCNNs and a pose estimator. Features are merged through a Temporal Feature Aggregation Module with self-attention and 1D convolution.
@@
NLA-SLR [214] targets semantically similar signs using a multistream backbone (RGB and pose heatmaps via S3D), followed by a head with language-aware label smoothing and an auxiliary inter-modality mix-up classifier. MSNN [116] uses five I3D streams (RGB and optical flow) plus ST-GCN for pose. hDNN-SLR [142] introduces a hybrid deep network with a 3DCNN for spatial detection, an attention-based BiLSTM for temporal modeling, and a VAE-based module for feature discrimination. Finally, Uni-Sign [102] proposes a unified multimodal architecture with three pose encoders and an RGB encoder fused via an attention-based module. Features are passed through temporal encoders and a pretrained LLM. Training involves pose-only and joint RGB-pose pretraining followed by downstream fine-tuning on ISLR and CSLR tasks.
@@ 5.3.3 Human-Computer Interaction
Motivated by the difficulty of recognizing fine-grained gestures in HCI applications, authors in [200] proposed HandSense, a method relying on depth and RGB data. Two 3DCNNs are deployed to extract features for each one modality. These features are fused and fed into a SVM to obtain the final prediction. Authors in [87] proposed an approach relying on RGBD Dynamic Images (RDDIs). These images, are obtained by the combination of RGB and estimated depth data. The resulted RDDI is fed into a modified Xception, resulting from removing the top 4 Xception modules, in order to extract meaningful features. These features are finally used to perform the classification.
@@ 5.3.4 Other
Inspired by the success of transformers, GestFormer [51] introduces a lightweight transformer-based framework for gesture-driven HRI. A 2DCNN extracts per-frame features, processed by PoolFormer layers (replacing attention) and two custom blocks for multiscale pooling/WT and token mixing. Each modality is processed independently, and outputs are averaged. In [49], a hybrid 3DCNN/LSTM model fuses RGB, depth, and skeleton data (from an enhanced pose estimator) via weighted sum. The fused input feeds a spatiotemporal model combining 2D/3D convolutions and ConvLSTM layers.
5.4 State-of-the-art Methods for Isolated Dynamic HGR
Recent IDHGR approaches demonstrate high architectural diversity. Pose-based methods often rely on spatial GCNs and temporal transformers [68, 71, 202, 204, 203, 165], or on ST-GCN variants [107, 79, 167]. Due to their limited standalone performance—mainly caused by pose estimation failures in challenging conditions (e.g., motion blur) [203]—pose-based models are frequently fused with simple RGB-based branches (e.g., I3D), using late fusion via score averaging [68, 71, 202, 204, 203, 104].
Appearance-based methods typically follow a multi-stream 3DCNN design to process individual modalities (e.g., RGB, depth, optical flow) [180, 206, 60, 100, 119, 195], employing late or multi-level fusion. Notably, SAM-SLR [82] and SAM-SLRv2 [81] show that adding more modalities—even with simple backbones—can significantly improve accuracy.
Advanced learning strategies further improve performance. Self-supervised learning (SSL) via masked pose reconstruction [202, 68, 71, 79, 203] enhances representation quality by injecting hand topology priors. Neural Architecture Search (NAS) has been used to discover optimal network structures [206, 195], while multi-task learning, as in Uni-Sign [102], boosts performance by leveraging auxiliary tasks and larger training corpora. Performance comparisons across benchmarks are summarized in Tables 4, 5, and 6.
Table 4: Brief description and reported accuracy of state-of-the-art methods on IsoGD [179].
Method Modality Architecture Description Training Setting Accuracy
3SCKI [180] RGBD, OF Three-stream 3DCNN with mid-level fusion Knowledge distillation between two instances of the pipeline, CLIP 65.87%
SlowFast+NAS [195] RGBD, OF Multi-stream and multi-rate 3DCNN with 3D-CDC and multi-level fusion NAS 66.23%
RAAR3DNet [206] RGBD Dual stream I3D with adoptable cells, spatial and temporal attention Multi-task loss and NAS 66.62%
De+Recouple [207] RGBD Dual-stream spatial + temporal with AE for multi-level fusion and mechanisms to reduce redundancy Self-distillation for inter-frame learning, extra supervision to enhance the AE 66.79%
CTFB [60] RGBD Dual-stream I3D with multi-level fusion - 69.02%
[100] RGBD Dual-stream Res3D-18 + average fusion Memory network to enhance the discriminative capability of the extracted features and information-theoretic loss function 74.08%
Table 5: Performance of uni-modal pose-based methods on SHREC [166] and DHG [37].
Method Architecture Split SHREC DHG
[120] Three-stream approach: attention-based spatial-temporal model, attention-based temporal-spatial model and MLP 14 97.01% 92.00%
28 92.78% 88.78%
MS-ISTGCN [167] Three-stream model: ST-GCN with adoptive spatial GC, extended temporal GC and attention 14 96.7% 93.7%
28 94.9% 91.2%
TD-GCN [107] ST-GCN variant with channel-dependent and temporal-dependent adjacency matrices 14 97.02% 93.9%
28 95.36% 91.4%
[101] Two-stream approach consisting of an attention-based dual branch module and MAE 14 96.9% 94.21%
28 94.17% 92.11%
SBI-DHGR [129] Multi-stream 1DCNN + BiLSTM 14 97% 94.64%
28 92.36% 91.79%
STST [201] Two-stage transformer-based approach 14 97.62% 94.82%
28 95.83% 93.18%
DSTSA-GCN [30] Three stacked stages, each one consisting of group channel-wise and temporal-wise GCN, followed by multiscale TCN 14 97.74% 95.04%
28 95.37% 93.57%
DRL-DDNet [46] DD-Net for classification and FS-Net for keyframe extraction 14 96.5% -
28 93.8% -
Table 6: Performance of the current SOTA methods on ISLR. PI, PC, OF and DF stands for per-instance accuracy, per-class accuracy, optical flow and depth flow, respectively. Methods marked with star (*) use additional training data.
Method Modality Architecture Description Training Setting Dataset PI PC
BEST [202] RGB, Pose d-VAE + GCN + Transformer & RGB-based encoder + late fusion SSL pretraining (masked tokens reconstruction) MSASL 71.21% 68.24%
WLASL 54.59% 52.12%
CSL500 97.7% -
SignBERT [68] RGB, Pose Spatial GCN, temporal and chilarity embedding + Transformer & RGB-based stream + late fusion SSL pretraining (MANO-based masked hand reconstruction) MSASL 71.24% 67.96%
WLASL 54.69% 52.08%
CSL500 97.6% -
SignBERT+ [71] RGB, Pose Extension of SignBERT with Spatial-Temporal Position Encoding and enhanced late fusion technique The same as SignBERT MSASL 73.71% 70.77%
WLASL 55.59% 53.33%
CSL500 97.8% -
MASA [203] RGB, Pose Spatial GCN + Transformer & RGB-based stream SSL pretraining (masked input reconstruction), contrastive learning WLASL 55.77% 53.13%
TMS-Net [38] Pose Multi-stream decoupled GCN with multi-level fusion - WLASL 56.4% -
SKIM [104] Pose Multi-stream channel-wise multi-scale GCN + late fusion Hand part mixing for data augmentation WLASL 57.73% 55.40%
[204] RGB, Pose GCN+Transformer and I3D + late fusion Contrastive learning to enhance invariance, consistency constraints between features of hand and trunk parts and between joints and motions MSASL 74.18% 72.18%
WLASL 58.06% 55.66%
CSL500 97.8% -
SAM-SLR-v2 [81] RGBD, Pose, HHA, OF, DF Extension of SAM-SLR-v2 with 3D graph stream and learnable weights for late fusion - WLASL 59.39% 56.63%
CSL500 99.00% -
NLA-SLR [214] RGB, Pose Four-stream network based on S3D with two frame rates and multi-level fusion Language-aware label Smoothing and inter-modality mixup with an auxiliary classifier MSASL 73.80% 70.95%
WLASL 61.26% 58.31%
Uni-Sign [102]* RGB, Pose A pretrained LLM process features extracted by pose, visual and temporal encoders Three-stage training, multi-task learning with additional training data MSASL 78.16% 76.97%
WLASL 63.52% 61.32%

### v2

@@ 5.2.2 Spatial + Temporal
Many models use a two-step approach, where the spatial domain is first processed by a spatial perception module to extract spatial features. These features are then passed into a sequential model for temporal aggregation, resulting in a combined spatial-temporal representation of the input sequence. For instance, GCN-BERT (GCN-BERT) utilizes a two-layer GCN for spatial modeling, followed by BERT for temporal modeling.
Other general-purpose methods adopt a multi-stream, two-stage pose-based approach. In (SBI-DHGR), three streams, each containing three 1D convolutional layers, are combined with a parallel module of average pooling layers. The extracted features are then concatenated and input into an LSTM, followed by a fully connected layer and softmax for classification. Similarly, STr-GCN (STr-GCN) employs a two-stage model where a GCN spatial feature extractor is first applied to obtain intra-frame features and then passes the features through a transformer encoder for modeling inter-frame temporal relationships. The final classification scores are generated by an MLP.
@@ 5.2.3 Short-Term Spatiotemporal + Long-Term Temporal
For example, gesture recognition in autonomous driving has become an important research area. In (MG-GCT), the authors propose a pose-based approach called MG-GCT for recognizing police traffic gestures. The model uses two graph convolution streams: one to process skeleton joints and another to analyze joint motions. The features extracted from both streams are fused and passed through a transformer encoder for temporal modeling, with the final classification performed by an MLP. In another approach, TD-GCN (TD-GCN) introduces a pose-based method for IDHGR. Unlike traditional methods that use fixed skeleton topologies, TD-GCN employs temporal-sensitive topology learning to enhance feature quality. First, graph convolution is used to extract high-level spatiotemporal representations, which are then used to calculate temporal and channel-independent adjacency matrices. These matrices help model the skeleton, and the features are fused using the proposed TCF module. Authors in (ISLR-STGCN-attention-LSTM-ozdemir2025cross) proposed a dual-stream backbone processing both hand and whole body keypoints using ST-GCN models. Cross-attention is then applied to focus on the most important features, which are subsequently pass through a LSTM to model long-term dependencies.
@@ 5.2.4 Temporal
In (TELEOPERATED_SURGICAL_ROBOT), the authors proposed a pose-based method for surgical robot teleoperation. Preprocessed skeleton joints are fed into an LSTM, and the resulting features are passed to an MLP for final classification. The model was evaluated on a custom dataset. Similarly, in (Sabater_alberto), a cross-domain gesture recognition approach was developed using geometric features (e.g., bone angles), which are processed by a TCN. The resulting features are summarized into a sequence descriptor, and a linear classifier predicts the final class.
To address data scarcity, the authors in (ADVERSARIAL_DATA_AUGM) introduced AVSN, an adversarial learning-based data augmentation framework. The method jointly optimizes augmentation parameters and the HGR classifier. A BiLSTM serves as the discriminator, while an MLP acts as the generator. The generator produces challenging augmented data that exploits weaknesses in the discriminator, which in turn learns to correctly classify gestures.
@@ 5.2.5 Other
Several works explore alternative architectures or skeleton representations beyond the main categories. In (FS-EFCNN), a Fourier CNN with BiLSTM extracts spatiotemporal features from 3D joints for ISLR, using a trainable binary mask for feature selection. Similarly, (SCT-DCNN) applies feature thresholding based on variance and importance, followed by deep CNN classification. To overcome color-coded map limitations, (MRCNNSA) introduces the Joint Motion Affinity Map (JMAM) with a multi-stream attention-based 2DCNN for feature extraction and recognition.
Key frame selection is addressed in (KEYFRAMES_DRL_SKELETON) via FS-Net, a reinforcement learning agent that selects informative frames for DD-Net classification. Dual-stream architectures such as STGCN-LSTM (STGCN-LSTM-10965662) combine ST-GCN and self-attention ConvLSTM for capturing global and local features, while (Yangke_patcog) proposes a lightweight variant merging attention-based and autoencoder streams. Finally, the Hierarchical Self-Attention Network (HAN) (HAN-liu2025han) employs modular attention across finger joints (J-Att, F-Att, T-Att, Fusion-Att) for efficient spatial–temporal modeling, achieving strong accuracy with just 0.05 GFLOPs.
@@ 5.3.1 General Purpose
In (Yunan_Li), a dual-stream 3DCNN with average fusion was proposed for RGBD gesture recognition, using a memory bank to retain discriminative features ($z$) while filtering redundant ones ($\upsilon$) by maximizing mutual information with class labels. Similarly, (RAAR3DNet) employed Neural Architecture Search (NAS) with Dynamic Static Attention (DSA) modules and a modified I3D backbone to emphasize motion and hand regions, supported by an auxiliary HeatmapNet and late RGBD fusion.
To better exploit modality interactions, (Benjia_Zhou) introduced a decoupling–recoupling framework combining independent spatial and temporal modeling with mid-level fusion through a shared MLP-AE module. Multi-level fusion was further explored in (CTFB_Hampiholi), integrating 3DCNNs with a Convolutional Transformer Fusion Block (CTFB) that fuses intermediate cross-stream features.
To address data scarcity, several works explored zero-shot learning (ZSL) by incorporating textual gesture descriptions encoded by pretrained language models (mainly BERT). In (ZS-GR), a dual-stream model combined a ViT for spatial modeling and an LSTM-MLP for temporal modeling, learning visual features aligned with textual embeddings. Similarly, (ZS-DHGR) fused skeleton-derived geometric features with RGB-C3D-LSTM-AE representations into a shared semantic space for classification. To reduce computational cost, (ConvMixFormer) replaced the transformer’s self-attention mechanism with convolutional layers and a gated feedforward network, resulting in the lightweight ConvMixFormer architecture.
Knowledge distillation and contrastive learning were applied in (3STREAM_DISTILLATION), where three 3DCNNs (RGB, Depth, Optical Flow) were fused with Coordinate Attention, using I3D as teacher and ResNet as student, while integrating CLIP for zero-shot recognition. Finally, (MFST) and (3D_CDC_NAS_TIP_2021) replaced standard convolutions with 3D Central Difference Convolution (3D-CDC), combining hybrid residual blocks, transformers, and NAS-based optimization for enhanced spatiotemporal modeling.
@@ 5.3.2 Sign Language Understanding
Recent works on Isolated Sign Language Recognition (ISLR) have focused on multimodal, pose-centric architectures combining RGB and skeleton data. SignBERT (SignBERT_ICCV) employs a spatial GCN with temporal and hand chirality encoding, feeding features into a transformer encoder and a hand-model-aware decoder for self-supervised pretraining, with parallel RGB predictions fused at inference. Its enhanced variant, SignBERT+ (SignBERT+), adds spatial-temporal position encoding and an improved RGB–pose fusion module (Fig. 6). Similarly, MASA (MASA) uses the same GCN+Transformer backbone with contrastive learning on pose reconstruction, while BEST (BEST) embeds pose triplets with BERT and applies late fusion with RGB data.
@@
Figure 6: The pose-based stream of SignBERT+ (SignBERT+), an extension of SignBERT (SignBERT_ICCV), first uses a GCN to extract gesture features, which are combined with spatial-temporal position encoding and processed by a transformer encoder. During self-supervised pretraining, a MANO-based decoder reconstructs masked inputs, and is replaced by a prediction head for downstream fine-tuning.
@@
To improve feature invariance, (Weichao-Zhao) applies a contrastive approach enforcing consistency among hand, trunk, and motion features, using the same GCN+Transformer backbone with late fusion of pose and RGB predictions. SAM-SLR (SAM_SLR) fuses six streams (RGB, depth, skeleton, optical flow, pose heatmaps, HHA), while SAM-SLR-v2 (jiang2021sign-sam-slr-v2) adds a 3D hand graph stream and learnable fusion weights.
VTN-HC/PF (VTN-HC-PF) processes hand-cropped images via a 2DCNN, transformer, and MLP for final predictions, with PF incorporating pose-based geometrical features. Motion History Images are used in (MOTION_HISTORY_Mercanoglu), analyzed by a 3DCNN with attention, and in a dual-stream setup combined with 2DCNN RGB features via weighted summation.
@@
To tackle data scarcity, ZSSLR (ZSSLR) uses a CLIP-based zero-shot framework, while (FSSLR) proposes a few-shot cross-lingual method combining two 3DCNNs and a pose estimator with a Temporal Feature Aggregation Module. NLA-SLR (NLA_SLR) targets semantically similar signs with a multistream S3D backbone and language-aware label smoothing. MSNN (MSNN) uses five I3D streams plus ST-GCN, and hDNN-SLR (hDNN-SLR) combines a 3DCNN, attention-based BiLSTM, and a VAE-based module. Finally, Uni-Sign (UNISIGNli2025) proposes a unified multimodal architecture with three pose encoders and an RGB encoder fused via attention, followed by temporal encoding and LLM-based pretraining for ISLR and CSLR.
@@ 5.3.3 Human-Computer Interaction
Motivated by the difficulty of recognizing fine-grained gestures in HCI applications, authors in (HandSense) proposed HandSense, a method relying on depth and RGB data. Two 3DCNNs are deployed to extract features for each one modality. These features are fused and fed into a SVM to obtain the final prediction. Authors in (mXception_and_RDDI) proposed an approach relying on RGBD Dynamic Images (RDDIs). These images, are obtained by the combination of RGB and estimated depth data. The resulted RDDI is fed into a modified Xception, resulting from removing the top 4 Xception modules, in order to extract meaningful features. These features are finally used to perform the classification.
@@ 5.3.4 Other
Inspired by the success of transformers, GestFormer (GestFormer) introduces a lightweight transformer-based framework for gesture-driven HRI. A 2DCNN extracts per-frame features, processed by PoolFormer layers (replacing attention) and two custom blocks for multiscale pooling/WT and token mixing. Each modality is processed independently, and outputs are averaged. In (DHGR_3D_POSE_HRI), a hybrid 3DCNN/LSTM model fuses RGB, depth, and skeleton data (from an enhanced pose estimator) via weighted sum. The fused input feeds a spatiotemporal model combining 2D/3D convolutions and ConvLSTM layers. Authors in (MDSI-fang2025mdsi) emphasize on two major issues, information redundancy and the difficulty to recognize similar gestures. Their approach leverages a CLIP encoder to provide qualitative embeddings and enhance the feature extractor through label smoothing as well as a semantic filter to incorporate multimodal data.
5.4 Performance of IDHGR Methods
The performance of IDHGR methods in the most used benchmarks (IsoGD (chalearn-lap), SHREC (SHREC), DHG (DHG-CVPRW), NVIDIA Hand Gesture or nvGesture (NVIDIA_GESTURE_R3DCNN_Molchanov), MSASL (MS-ASL), WLASL (WLASL), CSL500 (CSL500) and AUTSL (AUTSL)) is shown Table 3. For ISLR, Uni-Sign (UNISIGNli2025) surpasses relevant methods in terms of recognition accuracy, but requires powerful hardware to be executed fast, due to its large size (592.1M parameters, most of them contained on the language model). Thus, smaller but less accurate models like NLA-SLR (NLA_SLR), SignBERT+ (SignBERT_ICCV) and (Weichao-Zhao) are more reasonable solutions for real-world applications. Even though the multimodal SAM-SLR (SAM_SLR) and SAM-SLR-v2 (jiang2021sign-sam-slr-v2) are very accurate models for ISLR, they require a large number of modalities that are not fused effectively (using late fusion), which may be suboptimal. Pose-based methods for HCI and HRI including TD-GCN (TD-GCN), STST (STST) and DSTSA-GCN (DSTSA-CGN_cui2025dstsa) achieve high recognition rates with relatively low computational cost. Although, the computational complexity of pose estimation and the effect of its incorrect predictions in more challenging environments (e.g. fast movements) may reduce their overall performance. Appearance-based methods are not affected by off-the-shelf pose estimators but they comprise many more parameters (up to 251.4M (3D_CDC_NAS_TIP_2021)).
Table 3: Performance of IDHGR methods on various datasets. Notably the accuracy depends on the datasets’ difficulty; for example, methods with accuracy near to 100% on the AUTSL, do not perform the same way on more challenging datasets with thousands of classes (e.g. MSASL and WLASL).
SHREC DHG MSASL1000 WLASL2000
Method Year Params IsoGD 14 28 14 28 nvGesture PI PC PI PC CSL500 AUTSL
STGCN-LSTM (STGCN-LSTM-10965662) 2025 - - - - - - - - - - - 95.2% -
(ISLR-STGCN-attention-LSTM-ozdemir2025cross) 2025 - - - - - - - - - - - - 86.43%
HAN-2S (HAN-liu2025han) 2025 0.94M - 95% 92.86% 92.71% 89.15% - - - - - - -
DSTSA-GCN (DSTSA-CGN_cui2025dstsa) 2025 8.08M - 97.74% 95.37% 95.04% 93.57% - - - - - - -
Uni-Sign (UNISIGNli2025) 2025 592.1M - - - - - - 78.16% 76.97% 63.52% 61.32% - -
TMS-Net (TMS-NET) 2024 - - - - - - - - - 56.4% - - -
MSNN (MSNN) 2024 - - - - - - - 50.59% - 47.26% - - -
SKIM (SKIM) 2024 - - - - - - - - - 57.73% 55.40% - -
MASA (MASA) 2024 - - - - - - - - - 55.77% 53.13% - -
DSTA-SLR (DSTA_SLR_hu2024dynamic) 2024 7.4M - - - - - - 65.74% 62.31% 53.68% 51.17% 98.1%5 -
(Weichao-Zhao) 2024 - - - - - - - 74.18% 72.18% 58.06% 55.66% 97.8% -
TD-GCN (TD-GCN) 2024 1.36M - 97.02% 95.36% 93.9% 91.4% - - - - - - -
STST (STST) 2024 - - 97.62% 95.83% 94.82% 93.18% - - - - - - -
(Tianyu_Liu) 2024 - - - - - - - - - - - 98.87% 92.53%
SwC GR-MMixer (SwC-GR-MMixer) 2024 - - - - - - - - - - - 98.54% -
HWGAT (FDMSE-hierarchicalwindowedgraphattention) 2024 - - - - - - - - - 48.49% - - 95.80%
(Yangke_patcog) 2024 0.21M - 96.9% 94.17% 94.21% 92.11% - - - - - - -
(omar_ikne) 2024 - - 94.1% 90.0% - - - - - - - - -
BEST (BEST) 2023 - - - - - - - 71.21% 68.24% 54.59% 52.12% 97.7% -
SignBERT+ (SignBERT+) 2023 - - - - - - - 73.71% 70.77% 55.59% 53.33% 97.8% -
NLA-SLR (NLA_SLR) 2023 - - - - - - - 73.80% 70.95% 61.26% 58.31% - -
3SCKI (3STREAM_DISTILLATION) 2023 - 65.87% - - - - - - - - - - -
(Yunan_Li) 2023 - 74.08% - - - - - - - - - - -
CTFB (CTFB_Hampiholi) 2023 27.2M-37.1M 69.02% - - - - 87.76% - - - - - -
L-TSSI-DenseNet (L-TSSI-DenseNet) 2023 7.2M - - - - - - - - - - - 93.13%
TS3C-Net (TS3C-Net) 2023 28.7M - - - - - 87.55% - - - - - -
(Musa_IEEE_ACCESS) 2023 - - 97.01% 92.78% 92.00% 88.78% - - - - - - -
SBI-DHGR (SBI-DHGR) 2023 28M - 97% 92.36% 94.64% 91.79% - - - - - - -
STr-GCN (STr-GCN) 2023 - - 93.39% 89.20% - - - - - - - - -
DRL-DDNet (KEYFRAMES_DRL_SKELETON) 2023 - - 96.5% 93.8% - - - - - - - - -
De+Recouple (Benjia_Zhou) 2022 - 66.79% - - - - 91.70% - - - - - -
MS-ISTGCN (JOINTS_MOTION_VECTORS_Improved_ST-GCN) 2022 7.1M - 96.7% 94.9% 93.7% 91.2% - - - - - - -
SignBERT (SignBERT_ICCV) 2021 - - - - - - - 71.24% 67.96% 54.69% 52.08% 97.6% -
SlowFast+NAS+CDC (3D_CDC_NAS_TIP_2021) 2021 251.4M 66.23% - - - - - - - - - - -
RAAR3DNet (RAAR3DNet) 2021 - 66.62% - - - - 88.59% - - - - - -
SeST (SeST) 2021 - 60.27% - - - - - - - - - - -
SAM-SLR-v2 (jiang2021sign-sam-slr-v2) 2021 - - - - - - - - - 59.39% 56.63% 99.00% 98.53%
SAM-SLR (SAM_SLR) 2021 19.2M - - - - - - - - 58.73% 55.93% - 98.53%
(Hu_Zhou_Li_2021) 2021 - - - - - - - 69.39% 66.54% 51.39% 48.75% 98.3% -